A truck is moving at a speed of 36km/he is stopped by applying brakes which produce a uniform acceleration of -0.5m/s².How much distance will be covered by the scooter before it stops?


1
Expert's answer
2021-06-28T17:16:06-0400

We can find the distance covered by the scooter before it stops from the kinematic equation:


"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{0-(10\ \dfrac{m}{s})^2}{2\cdot(-0.5\ \dfrac{m}{s^2})}=100\ m."